# Learning Dynamic Bayesian Networks from Data: Foundations, First Principles and Numerical Comparisons

### v1

@@ 7.1 Model Validation Accuracy
For validation of the accuracy, we split the time series so that the first $70\,\%$ are used for training, and the remaining $30\,\%$ are used for testing. Then, we use the dbnR package to evaluate the log-likelihood of the train data given the predicted model. Results are presented in Tables 2 and 3.
Table 2: Log-likelihood for the favorable regime. TL indicates a setting that did not finish within the time limit, and E indicates a setting that ended in an error.
(3,30,10) (5,50,50) (10,100,200)
GFN -3.872058 144.9295 TL
Dynotears -7.249293 -68.31561 -1911.473
Gobnilp -8.655779 -37.15421 -1474.687
MCMC -6.983337 -71.15038 148.7103
PCMCI+ -9.333284 -144.704 -1986.396
MMHC -7.827745 -61.64009 TL
One Shot F. ILP -1.554613 31.7567 TL
Table 3: Log-likelihood for the high dimensional regime. TL indicates a setting that did not finish within the time limit, and E indicates a setting that ended in an error.
(3,5,10) (5,10,20) (10,20,40) (20,40,50) (30,60,100)
GFN -Inf 192.0897 TL TL TL
Dynotears 23.58121 35.43635 126.4152 -876.4649 -2071.534
Gobnilp 24.81038 44.50767 229.5804 TL TL
MCMC 32.12857 46.05138 208.4295 -457.6892 135.5351
PCMCI+ 23.6628 13.21597 36.11931 -1021.048 -2506.6
MMHC E E E E E
One Shot F. ILP 29.60606 75.58703 TL TL TL
@@ 7.2 Structure Identification
To evaluate the qualitative measures of the predicted structure, we compared the predictions with the ground truth adjacency matrix. The comparison was made using the structural Hamming distance, which is informally the number of edges that need to be either removed from or added to the predicted structural graph. The second measure is the AUROC, a standard metric that measures the area under the receiver operator characteristic. The results can be found in Tables 4 and 5.
Table 4: Expected SHD and AUROC for favorable dimensional regime for identification. TL indicates a setting that did not finish within the time limit, and E indicates a setting that ended in an error.
(3,30,10) (5,50,50) (10,100,200)
SHD AUROC SHD AUROC SHD AUROC
GFN 49.0 0.766 1030.0 0.871 16432.0 0.787
Dynotears 13.0 0.658 447.0 0.608 5578.0 0.583
Gobnilp 25.0 0.493 655.0 0.548 5018.0 0.562
MCMC 31.0 0.486 378.0 0.555 4321.0 0.687
PCMCI+ 19.0 0.5 242.0 0.610 5016.0 0.541
MMHC 20.0 0.650 516.0 0.552 TL TL
One Shot F. ILP 46.0 0.770 620.0 0.824 TL TL
Table 5: Expected SHD and AUROC for high dimensional regime. TL indicates a setting that did not finish within the time limit, and E indicates a setting that ended in an error.
(3,5,10) (5,10,20) (10,20,40) (20,40,50) (30,60,100)
SHD AUROC SHD AUROC SHD AUROC SHD AUROC SHD AUROC
GFN 57.212 0.797 442.0 0.728 3636.0 0.850 19228.0 0.863 90471 0.845
Dynotears 13.0 0.658 174.0 0.603 1126.0 0.558 2863.0 0.525 9384.0 0.528
Gobnilp 33.0 0.634 175.0 0.712 1022.0 0.663 TL TL TL TL
MCMC 43.0 0.647 228.0 0.545 904.0 0.664 2753.0 0.615 9801.0 0.614
PCMCI+ 19.0 0.5 121.0 0.5 822.0 0.519 2408.0 0.508 7639.0 0.534
MMHC E E E E E E E E E E
One Shot F. ILP 44.0 0.470 322.0 0.692 TL TL TL TL TL TL
8 Discussion and Conclusion
We hope this paper has provided a useful guide to the main principles behind learning the structure and parameters of a DBN. We focused on the fundamentals for the most simple cases, while targeting breadth in the scope of the various methodological approaches to learning these models from data.
There is an important aspect to DBNs that we did not discuss, as for the simple cases of learning it can be considered an orthogonal topic. This would be inference. DBNs are a generative model, so by themselves they do not accomplish any particular statistical decision test. However, one can perform various inference inqueries, such as the probability an instance of $X(2)$ with $X(1)=3.2$ and $Z=3$ be greater than $2.1$. One natural one for DBNs is a forward time forecast. Causal inference can also be performed through queries DBN models. Inference and approximate inference have a number of different procedures available, as far as efficiently and effectively sampling from the network.
Furthermore, inference algorithms are required in order to further extend DBN modeling to many real world datasets. For one, they become necessary for the expectation step in an Expectation-Maximization algorithm to learn structure with hidden variables. Often, with systems wherein the mechanism of action isn’t observed, a latent variable structure is able to model a rough set of possible dependencies that fits the observed data directed to and from in the graph.
For larger dimensions, IP approaches become computationally infeasible. In such a circumstance, given the Sample Complexity discussed in Section 3.
When data is plentiful, that is, millions and possibly easily available streaming samples, then neural network approaches can be effective. This suggests, for instance, the potential scalability of Generative Flow Networks [2] for instance. Reinforcement Learning is another common approach [74]. Otherwise, in the high-dimensional regime, wherein samples are finite but there are many covariates, Bayesian methods [6] or meta-heuristics are typically applied [33].
Acknowledgements
The authors would like to thank and Ondřej Kuželka for his suggestions and discussion on this work. This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101084642.

### v2

@@ 6.1 Model Validation Accuracy
%Test train split, perform some different inference tasks on the validation test set. %We evaluate the log-likelihood of the test set using dbnR package.
For accuracy’s validation, we split the time series so that the first \$70 %\$ are used for training, and the remaining \$30 %\$ are for testing. Then, we use the dbnR package to evaluate the log-likelihood of the test data given the predicted model. For Dynotears, only the structure was used to calculate the log-likelihood (so a binary adjacency matrix instead of the weighted one). Cycles were present in the Dynotears output networks so we used an algorithm that removed edges iteratively until cycles were no longer detected. Results are presented in Tables 6.1 and 6.1. For the error E in Table 6.1 for Gobnilp, a value -Inf was obtained for one of the datasets for that experiment ((3,5,10)). In the bnlearn package (from where dbnR imports the log-likelihood function), it is explained that logLik() returns -Inf if the data have probability or density equal to zero, which typically happens if the model is singular.
%See[pena2005learning]
%Favorable Data Regime
Table 2: Log-likelihood for the favorable regime. TL indicates a setting that did not finish within the time limit, and E indicates a setting that ended in an error, OOM: out of memory.
% % & (3,30,10)    &   (5,50,50)   &   (10,100,200)  &   (20,400,400)   &   (30,600,500)
[…]
GFN & -9.96\$±\$34.31 & -207.67\$±\$195.33 & -66740.91\$±\$22342.85 & -960522.9\$±\$46428.33 & -2542030\$±\$86678.95
Dynotears & -13.50\$±\$32.62 & -245.84\$±\$201.21 & -$6$1056631\$±\$39963.83 & -3369917\$±\$43506.22
Gobnilp & -6.89\$±\$33.65 & -188.74\$±\$205.84 & -51572.06\$±\$17188.01 & -865201.8\$±\$9170.71 & OOM
MCMC & -5.11\$±\$32.07 & -69.62\$±\$207.15 & -45776.51\$±\$15531.54 & -771035\$±\$786.78 & TL
PCMCI+ & -10.78\$±\$32.52 & -144.57\$±\$215.43 & -67263.58\$±\$23255.77 & TL & TL
HC-dbnR & -6.46\$±\$33.40 & -79.74\$±\$207.46 & -45782.72\$±\$15511.89 & TL & TL
% One Shot F. ILP &
%High Dimensional Regime
Table 3: Log-likelihood for the high dimensional regime. TL indicates a setting that did not finish within the time limit, and E indicates a setting that ended in an error, OOM: out of memory.
% & (3,5,10)    &   (5,10,20)   &   (10,20,40)  &   (20,40,50)  &   (30,60,100)
[…]
GFN & -36.89\$±\$9.98 & -341.29\$±\$17.37 & -2809.70\$±\$78.68 & -14166.43\$±\$247.86 & -61942.44\$±\$813.47
Dynotears & -32.07\$±\$7.94 & -329.96\$±\$17.44 & -2589.79\$±\$169.73 & -13860.83\$±\$233.30 & -66696.45\$±\$736.65
Gobnilp & E & -270.77\$±\$24.94 & -2151.40\$±\$76.75 & -10719.46\$±\$197.12 & OOM
MCMC & -29.89\$±\$8.82 & -237.45\$±\$15.95 & -1773.13\$±\$34.47 & -8755.98\$±\$127.74 & TL
PCMCI+ & -46.28\$±\$8.94 & -363.10\$±\$14.59 & -2898.88\$±\$80.93 & -14876.04\$±\$256.19 & TL
HC-dbnR & -31.04\$±\$11.34 & -240.75\$±\$17.61 & -1819.80\$±\$38.34 & -8991.89\$±\$157.52 & TL
% One Shot F. ILP & 29.60606 & 75.58703 & TL & TL & TL
We see that overall it appears that PCMCI+ maintains the highest accuracy, however, for larger problem sizes becomes impractical, at which point Dynotears and GFN become the two serious options.
@@ 6.2 Structure Identification
%Reconstruction distance to the correct graph %( Structural Hamming Distance)
To evaluate the qualitative measures of the predicted structure, we compared the predictions with the ground truth adjacency matrix. The comparison was made using the Structural Hamming distance (SHD), which is informally the number of edges that need to be either removed from or added to the predicted structural graph to convert it to the ground truth graph. The second measure is the AUROC, a standard metric that measures the area under the receiver operator characteristic. For GFN, we used a 99th percentile threshold on the posterior for each experiment to only retain the highest probabilities. The results can be found in Tables 6.2 and LABEL:table:hd_roc_shd. In PCMCI+ experiments for the high dimensional regime in Table LABEL:table:hd_roc_shd, the algorithm returns an empty matrix, which explains the 0.5 AUROC in all the experiments.
%Favorable Data Regime
Table 4: Expected SHD and AUROC for favorable dimensional regime for identification. TL indicates a setting that did not finish within the time limit, and E indicates a setting that ended in an error, OOM: out of memory.